### v1

@@ Abstract.
The aim of this paper is to study $m$-isometric weighted shifts with operator weights (both unilateral and bilateral). We obtain a characterization of such shifts by polynomials with operator coefficients. The procedure of construction of all $m$-isometric weighted shifts with positive weights is given. We answer the question when the weights of $m$-isometric shifts are commuting. The completion problem for $m$-isometric weighted shifts with operator weights is solved. We characterize $m$-isometric bilateral shifts the adjoints of which are also $m$-isometric. Several relevant examples are provided.
@@ 1. Introduction
$m$-isometric weighted shifts (with scalar weights) were characterized first byBermúdez et.al. in [8]. Some time later, in [1] there appeared a more convenient characterization of $m$-isometric shifts by real polynomials of degree at most $m-1$. Our goal is to generalize the results of [1]. The paper is organized as follows. In Section 2 we set up notation and terminology and review some of the standard facts on weighted shifts with operator weights. Section 3 is devoted to study $m$-isometric unilateral weighted shifts. Theorem 3.1 characterizes $m$-isometric unilateral weighted shifts with operator weights by polynomials of degree at most $m-1$ with operator coefficients; it may be viewed as a certain generalization of [1, Theorem 1]. In Theorem 3.4 we give the method of constructing $m$-isometric weighted shifts with positive and invertible weights. Theorem 3.10 answers the question when the weights of $m$-isometric unilateral weighted shifts commute; it turns out that they commute if and only if the coefficients of polynomial given by Theorem 3.1 commute. We also solve the completion problem for $m$-isometric unilateral weighted shifts with operator weights (see Theorem 3.9). Several examples are also provided. The aim of Section 4 is to provide counterparts of results presented in Section 3 for bilateral weighted shifts. Beyond that it is shown that the adjoint of $m$-isometric bilateral weighted shift with operator weights is also $m$-isometric if and only if the polynomial given by Theorem 4.4 is invertible (as an element of the algebra of polynomials with operator coefficients). This allows us to answer the question in [7, Problem 4.9] in negative.
@@ 2. Preliminaries
If $R$ is an algebra over a field $\mathbb{K}\in\left\{\mathbb{R},\mathbb{C}\right\}$, then $R[z]$ stands for the algebra of all polynomials with coefficients in $R$ with standard addition and multiplication. For $n\in\mathbb{N}$, $R_{n}[z]$ denotes the vector space over $\mathbb{K}$ of all polynomials of degree less than or equal to $n$.
@@
(2.1) $H=\mathcal{N}(T)\oplus\mathcal{R}(T^{\ast}).$
@@
$T$ is called positive if $\langle Tf,f\rangle>0$ for every $f\in H$. For $N\in\left\{\mathbb{N},\mathbb{Z}\right\}$ we define
@@
so $S^{\ast}(x_{j})_{j\in\mathbb{Z}}=(S_{j+1}^{\ast}x_{j+1})_{j\in\mathbb{Z}}$. Denote by $\widehat{S}$ the bilateral weighted shift with weights $(S_{-j+1}^{\ast})_{j\in\mathbb{Z}}$. Define $U\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ by the formula:
@@ 3. $m$-isometric unilateral weighted shifts
Let $H$ be a Hilbert space. Suppose $S\in\mathbf{B}(\ell^{2}(\mathbb{N}),H)$ is a unilateral weighted shifts with invertible weights $(S_{j})_{j\in\mathbb{Z}_{1}}$. For $m\in\mathbb{Z}_{1}$ the following conditions are equivalent:
@@
    Assume that the weights of $S$ are doubly commuting. Then $S$ is $m$-isometric if and only if there exists a polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ such that $p(0)=I$ and
@@
where $p_{j}\in\mathbb{R}_{m-1}[z]$ is such that $p_{j}(n)>0$ ($n\in\mathbb{N},j=1,\ldots,N$). We have
@@
Since $\mathcal{R}(S_{k}^{-2}\cdots S_{1}^{-1})=H$, (3.13) follows from the above. Now, again by [5, Lemma 2.1],
@@ 4. $m$-isometric bilateral weighted shifts
Define the sequence $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$ satisfying (4.1) and (4.2) by induction. Set $S_{1}=p(1)^{1/2}$ and $S_{0}=p(-1)^{-1/2}$. Clearly, $S_{1}$ and $S_{0}$ are positive and invertible and (4.1)-(4.2) hold for $n=-1,0,1$. Now assume that for some $k\in\mathbb{Z}_{1}$, $S_{-k+1},\ldots,S_{0},S_{1},\ldots,S_{k}$ are defined to be positive and invertible operators satisfying (4.1)–(4.2) for $n=-k,\ldots,k$. Then
$\widetilde{S}_{k+1}$ $=S_{1}^{-1}\cdots S_{k}^{-1}p(k+1)S_{k}^{-1}\cdots S_{1}^{-1},$
$\widetilde{S}_{-k}$ $=S_{-k+1}^{-1}\cdots S_{0}^{-1}p(-k-1)S_{0}^{-1}\cdots S_{-k+1}^{-1}$
are positive and invertible and so are $S_{k+1}=\widetilde{S}_{k+1}^{1/2}$ and $S_{-k}=\widetilde{S}_{-k}^{-1/2}$. Proceeding as in the proof of Theorem 3.4 we check that (4.1) holds for $n=k+1$. To verify that (4.2) holds for $n=-k-1$ observe that

### v2

@@ Abstract.
The aim of this paper is to study $m$-isometric weighted shifts with operator weights (both unilateral and bilateral). We obtain a characterization of such shifts by polynomials with operator coefficients. The procedure of construction of all $m$-isometric weighted shifts with positive weights is presented. We answer the question when the weights of $m$-isometric shift are commuting. The completion problem for $m$-isometric weighted shifts with operator weights is solved. We characterize $m$-isometric bilateral shifts the adjoints of which are also $m$-isometric.
@@ 1. Introduction
$m$-isometric weighted shifts (with scalar weights) were characterized first by Bermúdez et.al. in [8]. Some time later, in [1] there appeared a more convenient characterization of $m$-isometric shifts by real polynomials of degree at most $m-1$. Our goal is to generalize the results of [1]. The paper is organized as follows. In Section 2 we set up notation and terminology and review some of the standard facts on weighted shifts with operator weights. Section 3 is devoted to study $m$-isometric unilateral weighted shifts. Theorem 3.1 characterizes $m$-isometric unilateral weighted shifts with operator weights by polynomials of degree at most $m-1$ with operator coefficients; it may be viewed as a certain generalization of [1, Theorem 1]. In Theorem 3.4 we give the method of constructing all $m$-isometric weighted shifts with positive and invertible weights. Theorem 3.10 answers the question when the weights of $m$-isometric unilateral weighted shifts commute; it turns out that they commute if and only if the coefficients of polynomial given by Theorem 3.1 commute. We also solve the completion problem for $m$-isometric unilateral weighted shifts with operator weights (see Theorem 3.9). Several examples are also provided. The aim of Section 4 is to provide counterparts of results presented in Section 3 for bilateral weighted shifts. Beyond that it is shown that the adjoint of $m$-isometric bilateral weighted shift with operator weights is also $m$-isometric if and only if the polynomial given by Theorem 4.4 is invertible (as an element of the algebra of polynomials with operator coefficients). This allows us to answer the question in [7, Problem 4.9] in negative.
@@ 2. Preliminaries
If $R$ is an algebra over a field $\mathbb{K}\in\left\{\mathbb{R},\mathbb{C}\right\}$, then $R[z]$ stands for the algebra of all polynomials with coefficients in $R$ with standard addition and multiplication. For $n\in\mathbb{N}$, $R_{n}[z]$ denotes the vector space over $\mathbb{K}$ of all polynomials with coefficients in $R$ of degree less than or equal to $n$.
@@
(2.1) $H=\mathcal{N}(T)\oplus\overline{\mathcal{R}(T^{\ast})}.$
@@
$T$ is called positive if $\langle Tf,f\rangle>0$ for every $f\in H$, $f\not=0$. For $N\in\left\{\mathbb{N},\mathbb{Z}\right\}$ we define
@@
so $S^{\ast}(x_{j})_{j\in\mathbb{Z}}=(S_{j+1}^{\ast}x_{j+1})_{j\in\mathbb{Z}}$ for $(x_{j})_{j\in\mathbb{Z}}\in\ell^{2}(\mathbb{Z},H)$. Denote by $\widehat{S}$ the bilateral weighted shift with weights $(S_{-j+1}^{\ast})_{j\in\mathbb{Z}}$. Define $U\in\mathbf{B}(\ell^{2}(\mathbb{Z},H))$ by the formula:
@@ 3. $m$-isometric unilateral weighted shifts
Let $H$ be a Hilbert space. Suppose $S\in\mathbf{B}(\ell^{2}(\mathbb{N}),H)$ is a unilateral weighted shift with invertible weights $(S_{j})_{j\in\mathbb{Z}_{1}}$. For $m\in\mathbb{Z}_{1}$ the following conditions are equivalent:
@@
    Assume that the weights of $S$ are pairwise doubly commuting. Then $S$ is $m$-isometric if and only if there exists a polynomial $p\in\mathbf{B}(H)_{m-1}[z]$ such that $p(0)=I$ and
@@
where $p_{j}\in\mathbb{R}_{m-1}[z]$ are such that $p_{j}(n)>0$ ($n\in\mathbb{N},j=1,\ldots,N$). We have
@@
Since $\mathcal{R}(S_{k-1}^{-2}\cdots S_{1}^{-1})=H$, (3.13) follows from the above. Now, again by [5, Lemma 2.1],
@@ 4. $m$-isometric bilateral weighted shifts
Define the sequence $(S_{j})_{j\in\mathbb{Z}}\subset\mathbf{B}(H)$ satisfying (4.1) and (4.2) by induction. Set $S_{1}:=p(1)^{1/2}$ and $S_{0}:=p(-1)^{-1/2}$. Clearly, $S_{1}$ and $S_{0}$ are positive and invertible and (4.1)-(4.2) hold for $n=-1,0,1$. Now assume that for some $k\in\mathbb{Z}_{1}$, $S_{-k+1},\ldots,S_{0},S_{1},\ldots,S_{k}$ are defined to be positive and invertible operators satisfying (4.1)–(4.2) for $n=-k,\ldots,k$. Then
$\widetilde{S}_{k+1}$ $:=S_{1}^{-1}\cdots S_{k}^{-1}p(k+1)S_{k}^{-1}\cdots S_{1}^{-1},$
$\widetilde{S}_{-k}$ $:=S_{-k+1}^{-1}\cdots S_{0}^{-1}p(-k-1)S_{0}^{-1}\cdots S_{-k+1}^{-1}$
are positive and invertible and so are $S_{k+1}:=\widetilde{S}_{k+1}^{1/2}$ and $S_{-k}:=\widetilde{S}_{-k}^{-1/2}$. Proceeding as in the proof of Theorem 3.4 we check that (4.1) holds for $n=k+1$. To verify that (4.2) holds for $n=-k-1$ observe that